# PAC-Bayes Analysis for Recalibration in Classification

### v1

@@ A.6 Discussion about the assumption of Theorem 4
This is caused by the evaluation of the coefficient of the Lemma 1. In the derivation of Eq. (A.5), if we do not assume that $f_{w}$ has the probability density, then there is a possibility that the training data is allocated to the point at the exact grid points of the hypercube constructed by the bins. The grid point is the intersection of $2^{K}$ small regions, and thus, the $c_{i}$s of Lemma 1 becomes exponentially large. Owing to this, the coefficient of Eq. (A.5) is significantly larger compared with the setting when we assume that $f_{w}$ has the probability density, where we do not need to consider such a worst case. In this way, the order of the bias significantly improves by assuming that $f_{w}$ has the probability density.
Appendix B Proofs and discussion of Section 4
@@
B.1 Proofs for Corollary 1 and 2
When we recalibrate the model, we fix $w$. Thus under the $f_{w}$, the given recalibration data $S_{\mathrm{re}}=\{(X_{m},Y_{m})\}_{m=1}^{n_{\mathrm{re}}}$, it is transformed into $\tilde{S}_{\mathrm{re}}=\{(f_{w}(X_{m}),Y_{m})\}_{m=1}^{n_{\mathrm{re}}}$ and samples in $\tilde{S}_{\mathrm{re}}$ are i.i.d by definition. Then we apply our developed PAC-Bayes bounds in Appendix A for this new dataset.
As for Corollary 1, we can prove it exactly in the same way as shown in Appendix A.2.3. Simply replacing $S_{\mathrm{tr}}$ with $\tilde{S}_{\mathrm{re}}$ in Appendix A.2.3 and apply it to the standard derivation of PAC-Byaeisan bound shown in Eq. (39). This concludes the proof.
@@ B.1 Proofs for Corollary 1 and 2
$\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{V}(f_{w}),S_{\mathrm{re}},\mathrm{CE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda},$ (89)
$\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{V}(f_{w}),S_{\mathrm{re}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}.$ (90)
We can prove this exactly in the same way as shown in Appendix A.3. Simply replacing $S_{\mathrm{tr}}$ with $\tilde{S}_{\mathrm{re}}$ in Appendix A.3. This concludes the proof.
B.2 Discussion about the conditional KL divergence
@@
$\mathbb{E}_{\rho(W|s_{\mathrm{tr}})}\mathbb{E}_{\tilde{\rho}}\mathrm{gen}(\eta_{V}(f_{W}),S_{\mathrm{re}},\mathrm{ECE})\leq\frac{\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}[\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})]+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n}}{\lambda}.$ (91)
@@
The proof is almost the same as Appendix A.2.3. The difference is how we utilize the Chernoff-bounding technique. Here we only show the different parts. After obtaining the upper bound of the exponential moment $f(t,n_{\mathrm{re}})$ as in Eq. (52), we substitute this into Eq. (39), which can be written as follows; for any $w$ and $v$ we have
$\mathbb{E}_{S_{\mathrm{re}}}e^{t|\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\leq 1.$ (92)
@@
$\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\mathbb{E}_{\tilde{\pi}}\mathbb{E}_{S_{\mathrm{re}}}e^{t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\leq 1.$
@@
$\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\tilde{\pi}}\left[{\rm e}^{t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\right]\leq 1,$
@@
$\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\mathbb{E}_{\tilde{\pi}}\left[{\rm e}^{t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\right]\leq 1.$
@@
$\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\left[{\rm e}^{\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]\leq 1.$
@@
$\mathbb{E}_{S_{\mathrm{re}}}\left[{\rm e}^{\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]$
$\leq\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\left[{\rm e}^{\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]\leq 1.$ (93)
@@
$\mathbb{E}_{S_{\mathrm{re}}}\left[{\rm e}^{\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]$
$\leq\mathbb{E}_{S_{\mathrm{re}}}\left[{\rm e}^{\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{V}(f_{W}),S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]\leq 1.$ (94)
After this, we simply follow the derivation of the standard PAC-Bayes derivation in Appendix A.3. ∎
B.3 Discussion for Section 4.2
@@
$\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}|\hat{Y}_{C}-\eta_{v}(f_{w}(X))_{C}|^{2}\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}\|\hat{Y}-\eta_{v}(f_{w}(X))\|_{2}^{2}$ (95)
where $\hat{Y}_{C}$ is the $C$-th dimension of the one-hot encoding of the label $Y$ and the first inequality is followed by Hölder inequality and the Jensen inequality. As for the binary classification, by Hölder inequality and the Jensen inequality, we have
$\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}|Y-\eta_{v}(f_{w}(X))_{1}|^{2},$ (96)
since the recalibrated function is $\eta_{v}(f_{w}(X))\in\Delta^{2}$ and its 1st dimension represents the predicted probability of the label is $1$.
Then from the estimation bias bound in Eq. (14), we have
$\mathbb{E}_{\tilde{\rho}}\mathrm{TCE}(\eta_{v}(f_{w}))$
$\leq\mathbb{E}_{\tilde{\rho}}\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})+\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n_{\mathrm{re}}}}{\lambda}$
$\leq\mathbb{E}_{\tilde{\rho}}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}\|\hat{Y}-\eta_{V}(f_{w}(X))\|_{2}^{2}+\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n_{\mathrm{re}}}}{\lambda},$ (97)
@@
Next, we discuss the objective function of Eq. (16). This objective function can guarantee the performance of test accuracy and TCE simultaneously as follows; Assume that $l_{\mathrm{acc}}:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ and define the training loss as $\hat{L}_{\mathrm{acc}}(V,S_{\mathrm{tr}})\coloneqq\frac{1}{n_{\mathrm{tr}}}\sum_{(X,Y)\in S_{\mathrm{tr}}}^{n_{\mathrm{tr}}}l_{\mathrm{acc}}(Y,\eta_{v}(f_{w}(X)))$ and the (expected) test loss as $L_{\mathrm{acc}}(v,\mathcal{D})\coloneqq\mathbb{E}_{Z}l_{\mathrm{acc}}(Y,\eta_{v}(f_{w}(X)))$. Under this setting, we have
@@
$\mathbb{E}_{\tilde{\rho}}(L_{\mathrm{acc}}(V,\mathcal{D})+\mathrm{TCE}(\eta_{V}(f_{w}))$ $\leq\mathbb{E}_{\tilde{\rho}}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}(l_{\mathrm{acc}}(Y,\eta_{V}(f_{w}(X)))+\|\hat{Y}-\eta_{V}(f_{w}(X))\|_{2}^{2})$
$\quad+\frac{1+L}{B}+\frac{2\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\frac{33\lambda^{2}}{8n_{\mathrm{re}}}+2\log\frac{2}{\varepsilon}}{\lambda},$ (98)
@@
By considering the union bound with Eq. (99) and Eq. (B.3), setting $\varepsilon$ as $\varepsilon/2$ in each high probability bound, we obtain the result. ∎
@@
KITTI [8] $2$ $16000$ $1000$ $8000$
PCam [31] $2$ $22768$ $1000$ $9000$
MNIST [19] $10$ $60000$ $1000$ $9000$
CIFAR-100 [15] $100$ $50000$ $1000$ $9000$
Appendix C Details of experimental settings

### v2

@@ B.9 Discussion about the assumption of Theorem 3
This is caused by the evaluation of the coefficient of the Lemma 1. In the derivation of Eq. (B.8), if we do not assume that $f_{w}$ has the probability density, then there is a possibility that the training data is allocated to the point at the exact grid points of the hypercube constructed by the bins. The grid point is the intersection of $2^{K}$ small regions, and thus, the $c_{i}$s of Lemma 1 becomes exponentially large. Owing to this, the coefficient of Eq. (B.8) is significantly larger compared with the setting when we assume that $f_{w}$ has the probability density, where we do not need to consider such a worst case. In this way, the order of the bias significantly improves by assuming that $f_{w}$ has the probability density.
B.10 Discussion about the selected labels for the ECE
Here we consider the setting that we only focuses on $K^{\prime}\in[1,K]$ classes. Without loss of generality, we focus on the set of labels defined as $\{0,\dots,K^{\prime}-1\}$. We then define $f^{\prime}_{w}(x)=(f_{w}(x)_{1},\dots,f_{w}(x)_{K^{\prime}})$. We also regard $P(Y=0|f^{\prime}_{w}(X)),\dots,P(Y=K^{\prime}-1|f^{\prime}_{w}(X))$ as the $K^{\prime}$ dimensional vector given $w$ and $x$. We also define the one-hot vector $\mathbf{e}^{\prime}_{Y}\coloneqq((\mathbf{e}_{Y})_{1},\dots,(\mathbf{e}_{Y})_{K^{\prime}})$ where $(\mathbf{e}_{Y})_{k}$ is the $k$-the element of $\mathbf{e}_{Y}$.
Under these notations, we can define
$CE_{K^{\prime}}(f_{w})\coloneqq\mathbb{E}\|P(Y|f^{\prime}_{w}(X))-f^{\prime}_{w}(X)\|_{1},$ (110)
where $\|\cdot\|_{p}$ is the $L^{p}$ distance in $\mathbb{R}^{K^{\prime}}$. The $CE_{K^{\prime}}$ metric measures the calibration performance of models for only $K^{\prime}$ classes.
Let us consider the following binning scheme to estimate $CE_{K^{\prime}}$. Since we want to estimate the $K^{\prime}$-dimensional conditional probability in $[0,1]^{K^{\prime}}$ by binning, we split each dimension $[0,1]$ into $B^{\prime}$ bins of the same width. After doing so, there are $B=(B^{\prime})^{K^{\prime}}$ bins, and $[0,1]^{K^{\prime}}$ is split into $B$ small regions. We refer to these regions as $\mathcal{I}\coloneqq\{I_{i}\}_{i=1}^{B}$. In this case, the ECE of $\mathrm{CE}_{K}$ can be defined as
$\mathrm{ECE}_{K^{\prime}}(f_{w},S_{e})\coloneqq\sum_{i=1}^{B}p_{i}\|\bar{f}^{\prime}_{i,S_{e}}-\bar{p}^{\prime}_{i,S_{e}}\|_{1}$ (111)
where $|I_{i}|\coloneqq\sum_{m=1}^{n_{e}}\mathbbm{1}_{f^{\prime}_{w}(x_{m})\in I_{i}}$, $\hat{p}_{i}\coloneqq|I_{i}|/n_{e}$, $\bar{f}^{\prime}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f^{\prime}_{w}(x_{m})\in I_{i}}f^{\prime}_{w}(x_{m})$, and $\bar{p}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f^{\prime}_{w}(x_{m})\in I_{i}}\mathbf{e}^{\prime}_{y_{m}}$.
For these definitions, we can derive the following bias
$\mathrm{Bias}(f_{w},S_{\mathrm{te}},CE_{K^{\prime}})\coloneqq|CE_{K^{\prime}}(f_{w})-\mathrm{ECE}_{K^{\prime}}(f_{w},S_{\mathrm{te}})|$ (112)
and this can be analyzed exactly in the same way as Appendix B.8. Then we have $\mathrm{Bias}(f_{w},S_{\mathrm{te}},CE_{K^{\prime}})=\mathcal{O}(1/n_{\mathrm{te}}^{1/(K^{\prime}+2)})$ under $B=\mathcal{O}(n_{\mathrm{te}}^{\frac{1}{K^{\prime}+2}})$. We can obtain the similar generalization bound.
Appendix C Proofs and discussion of Section 4
@@
C.1 Proofs for Corollary 1 and 2
When we recalibrate the model, we fix $w$. Thus under the $f_{w}$, the given recalibration data $S_{\mathrm{re}}=\{(X_{m},Y_{m})\}_{m=1}^{n_{\mathrm{re}}}$, it is transformed into $\tilde{S}_{\mathrm{re}}=\{(f_{w}(X_{m}),Y_{m})\}_{m=1}^{n_{\mathrm{re}}}$ and samples in $\tilde{S}_{\mathrm{re}}$ are i.i.d by definition. Then we apply our developed PAC-Bayes bounds in Appendix B for this new dataset.
As for Corollary 1, we can prove it exactly in the same way as shown in Appendix B.6. Simply replacing $S_{\mathrm{tr}}$ with $\tilde{S}_{\mathrm{re}}$ in Appendix B.6 and apply it to the standard derivation of PAC-Byaeisan bound shown in Eq. (67). This concludes the proof.
@@ C.1 Proofs for Corollary 1 and 2
$\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{v}\circ f_{w},S_{\mathrm{re}},\mathrm{CE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda},$ (113)
$\mathbb{E}_{\tilde{\rho}}\mathrm{Bias}(\eta_{v}\circ f_{w},S_{\mathrm{re}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}.$ (114)
We can prove this exactly in the same way as shown in Appendix B.3. Simply replacing $S_{\mathrm{tr}}$ with $S_{\mathrm{re}}$ in Appendix B.3. This concludes the proof.
C.2 Discussion about the conditional KL divergence
@@
$\mathbb{E}_{\rho(W|s_{\mathrm{tr}})}\mathbb{E}_{\tilde{\rho}}\mathrm{gen}(\eta_{v}\circ f_{w},S_{\mathrm{re}},\mathrm{ECE})\leq\frac{\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}[\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})]+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n}}{\lambda}.$ (115)
@@
The proof is almost the same as Appendix B.6. The difference is how we utilize the Chernoff-bounding technique. Here we only show the different parts. After obtaining the upper bound of the exponential moment $f(t,n_{\mathrm{re}})$ as in Eq. (85), we substitute this into Eq. (67), which can be written as follows; for any $w$ and $v$ we have
$\mathbb{E}_{S_{\mathrm{re}}}e^{t|\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\leq 1.$ (116)
@@
$\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\mathbb{E}_{\tilde{\pi}}\mathbb{E}_{S_{\mathrm{re}}}e^{t|\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\leq 1.$
@@
$\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\tilde{\pi}}\left[{\rm e}^{t|\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\right]\leq 1,$
@@
$\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\mathbb{E}_{\tilde{\pi}}\left[{\rm e}^{t|\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|-f(t,n_{\mathrm{re}})}\right]\leq 1.$
@@
$\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\rho(w|s_{\mathrm{tr}})}\left[{\rm e}^{\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]\leq 1.$
@@
$\mathbb{E}_{S_{\mathrm{re}}}\left[{\rm e}^{\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]$
$\leq\mathbb{E}_{S_{\mathrm{re}}}\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\left[{\rm e}^{\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]\leq 1.$ (117)
@@
$\mathbb{E}_{S_{\mathrm{re}}}\left[{\rm e}^{\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]$
$\leq\mathbb{E}_{S_{\mathrm{re}}}\left[{\rm e}^{\mathbb{E}_{\rho(w|S_{\mathrm{tr}})}\sup_{\tilde{\rho}\in\mathcal{P}(\mathcal{W})}\mathbb{E}_{\tilde{\rho}}t|\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|-\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})-f(t,n_{\mathrm{re}})}\right]\leq 1.$ (118)
After this, we simply follow the derivation of the standard PAC-Bayes derivation in Appendix B.3. ∎
C.3 Discussion for Section 4.2
@@
$\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})^{2}\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}|(\mathbf{e}_{Y})_{C}-\eta_{V}\circ f_{w}(X)_{C}|^{2}\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}\|\mathbf{e}_{Y}-\eta_{V}\circ f_{w}(X)\|_{2}^{2}$ (119)
where $(\mathbf{e}_{Y})_{C}$ is the $C$-th dimension of the one-hot encoding of the label $Y$ and the first inequality is followed by Hölder inequality and the Jensen inequality. As for the binary classification, by Hölder inequality and the Jensen inequality, we have
$\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})^{2}\leq\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}|Y-\eta_{V}\circ f_{w}(X)_{1}|^{2},$ (120)
since the recalibrated function is $\eta_{V}\circ f_{w}(X)\in\Delta^{2}$ and its 1st dimension represents the predicted probability of the label is $1$.
Next we connect the square of the ECE to our derived PAC-Bayes bound; note that
$|\mathrm{TCE}(\eta_{v}\circ f_{w})^{2}-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})^{2}|$ (121)
$\leq|(\mathrm{TCE}(\eta_{v}\circ f_{w})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}}))(\mathrm{TCE}(\eta_{v}\circ f_{w})+\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}}))|$ (122)
$\leq 2|\mathrm{TCE}(\eta_{v}\circ f_{w})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|.$ (123)
Thus, taking the expectation over $v$, we have
$\frac{1}{2}\mathbb{E}_{\tilde{\rho}}|\mathrm{TCE}(\eta_{v}\circ f_{w})^{2}-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})^{2}|$ $\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n_{\mathrm{re}}}}{\lambda}$ (124)
Thus, combining the above results, we have
$\mathbb{E}_{\tilde{\rho}}\mathrm{TCE}(\eta_{v}\circ f_{w})^{2}$ $\leq\mathbb{E}_{\tilde{\rho}}\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})^{2}+2\frac{1+L}{B}+2\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n_{\mathrm{re}}}}{\lambda}$
$\leq\mathbb{E}_{\tilde{\rho}}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}\|\mathbf{e}_{Y}-\eta_{V}\circ f_{w}(X)\|_{2}^{2}+2\frac{1+L}{B}+2\frac{\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+B\log 2+\log\frac{1}{\varepsilon}+\frac{4\lambda^{2}}{n_{\mathrm{re}}}}{\lambda}$ (125)
@@
Next, we discuss the objective function of Eq. (26). This objective function can guarantee the performance of test accuracy and TCE simultaneously as follows; Assume that $l_{\mathrm{acc}}:\mathcal{Y}\times\mathcal{Y}\to[0,1]$ and define the training loss as $\hat{L}_{\mathrm{acc}}(V,S_{\mathrm{tr}})\coloneqq\frac{1}{n_{\mathrm{tr}}}\sum_{(X,Y)\in S_{\mathrm{tr}}}^{n_{\mathrm{tr}}}l_{\mathrm{acc}}(Y,\eta_{V}\circ f_{w}(X))$ and the (expected) test loss as $L_{\mathrm{acc}}(v,\mathcal{D})\coloneqq\mathbb{E}_{Z}l_{\mathrm{acc}}(Y,\eta_{V}\circ f_{w}(X))$. Under this setting, we have
@@
$\mathbb{E}_{\tilde{\rho}}(L_{\mathrm{acc}}(V,\mathcal{D})+\mathrm{TCE}(\eta_{v}\circ f_{w})^{2})$ $\leq\mathbb{E}_{\tilde{\rho}}\frac{1}{n_{\mathrm{re}}}\sum_{(X,Y)\in S_{\mathrm{re}}}^{n_{\mathrm{re}}}(l_{\mathrm{acc}}(Y,\eta_{V}\circ f_{w}(X))+\|\mathbf{e}_{Y}-\eta_{V}\circ f_{w}(X)\|_{2}^{2})$
$\quad+2\frac{1+L}{B}+\frac{3\mathrm{KL}(\tilde{\rho}\|\tilde{\pi})+2B\log 2+\frac{65\lambda^{2}}{8n_{\mathrm{re}}}+3\log\frac{2}{\varepsilon}}{\lambda},$ (126)
@@
By considering the union bound with Eq. (127) and Eq. (C.3), setting $\varepsilon$ as $\varepsilon/2$ in each high probability bound, we obtain the result. ∎
@@
KITTI [Geiger, 2012] $2$ $16000$ $1000$ $8000$
PCam [Veeling et al., 2018] $2$ $22768$ $1000$ $9000$
MNIST [LeCun et al., 1989] $10$ $60000$ $1000$ $9000$
CIFAR-100 [Krizhevsky, 2009] $100$ $50000$ $1000$ $9000$
Appendix D Details of experimental settings